• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kumbartzki

    Log entry time 15:59:34 on February15,2009

    Entry number 261322

    keyword=Day shift summary

    At start of shift data taking continues.
    Run 20205 / 1450
    08:25 Run 1451
    10:15 At the end of Run 1454 a NMR was done without stopping 20205
    We stopped that run minutes after the NMR sweep and started new.
    The BB alarms continue about 3/hr.
    We noticed periods of up to 10 minutes with very low beam current.
    The operators did not notice. Once a while there is a lock problem.
    We have to keep an eye on that. MCC claims to have fixed the problem for now.
    Smooth running for the rest of time
    13:33 Runs 20207 /1464